OpenStudy (anonymous):

-3 (5 -15n) = 15(3n - 2)

OpenStudy (mathmale):

Please include the instructions with each problem you post. Thanks.

OpenStudy (akashdeepdeb):

What is the question here @metalblade ?

OpenStudy (akashdeepdeb):

I am guessing, in this problem you jut have to find n. It is done like this: -3 (5 -15n) = 15(3n - 2) Use the distributive property to get rid of the parenthesis -3 * 5 - 3*(-15n) = 15*3n - 15*2 Simplify this now -15 + 45n = 45n - 30 We see that, 45n cancels out. Thus, this equation cannot be solved.

OpenStudy (anonymous):

-3(5-15n)=15(3n-2) -15+45n=45n-30 this problem does not work as you wrote it, basically -15 cannot equal-30 the problem has no solution as written

OpenStudy (anonymous):

C.S.Calvin is right. There is no solution.